Investing. Do Xuan Nam is going to invest $1000 at a simple interest rate of 4%. How long will it take for the investment to be worth %1040?
Answer by stanbon(75887)

Investing. Do Xuan Nam is going to invest $1000 at a simple interest rate of 4%. How long will it take for the investment to be worth %1040?
---------
A(t) = P(1+rt)
-------
1040 = 1000(1+0.04t)
-------
1.04 = 1+0.04t
0.04 = 0.04t
time = 1 year
=================
Cheers,
Stan H.
--------------